# Advanced Architectures Integrated with Agentic AI for Next-Generation Wireless Networks

### v2

@@ Abstract
Index Terms:
6G architectures, AI Agents, constrained AI, serverless computing, optical networks, deep learning, goal-oriented communication
@@ I Introduction
One of the key innovations of this paper lies in the design of simplified network architectures that separate the control and user planes, allowing operators to deploy and manage services more efficiently through AI Agents while seamlessly integrating new network domains. This approach also addresses the limitations of existing Service-Based Architectures (SBA) and provides a clear path for the long-term evolution of 6G networks. The increased flexibility of such architectures ensures that devices can operate as dynamic network nodes, enabling uninterrupted roaming across multiple operators and network technologies. In addition to architectural simplification, the integration of AI Agents Artificial Intelligence (AI) into network management [3] opens up new opportunities for optimization, especially in environments with energy constraints. Agentic AI-based networks can make intelligent decisions based on real-time data, but these models must work within strict energy and security constraints to ensure sustainable and secure operations. The concept of constrained AI operations, which is explored in this paper, is crucial for achieving these goals. It focuses on optimising energy consumption and learning in real-time while maintaining security protocols. In addition, the emergence of serverless computing and Function-as-a-Service (FaaS) offers new solutions for the orchestration of services in distributed network environments [11]. Serverless computing decouples application execution from the underlying infrastructure and enables real-time execution of lightweight functions at the network edge. This paradigm shift supports the flexible placement of AI agents and orchestration of tasks, a key requirement for the future 6G ecosystem where services must dynamically adapt to changing network conditions.
@@
Another fundamental component of future 6G architectures is the use of autonomous cognitive agents that introduce decentralized decision-making processes [15]. These agents work without a central controller and enable dynamic and spontaneous interactions that optimize network resources and operations. This is a significant departure from traditional static architectures and enables more adaptable and resilient networks that can meet future requirements. The concept of goal-oriented communication, a revolutionary approach that enables AI-equipped devices to transmit only the information necessary to achieve specific goals [12]. This reduces unnecessary communication, computing load and energy consumption, while allowing devices to communicate more efficiently. Such innovations are crucial for next-generation machine-to-machine (M2M) communication, where optimising resource usage is key. In addition, the potential for cloudification and management of RAN functions in space underscores the increasing role of distributed data centers, including those in space, in the dynamic orchestration of RAN functions [5]. This approach contrasts with existing static methods of resource allocation and provides a more flexible and adaptable solution that can meet the diverse requirements of 6G communication scenarios. Cloudification of RAN functions is expected to play a critical role in achieving ultra- reliable, low-latency communications for a wide range of applications. The concept of Neural Radio Protocol Stacks (NRPS) can be used to automate the generation of radio protocol stacks tailored to specific network requirements using Agentic AI techniques. NRPSs can be a significant advance in the development of customizable and efficient communication protocols. By integrating Agentic AI into the development of radio stacks [14], these solutions can offer a high degree of flexibility and adaptability, addressing issues such as interpretability, cross-vendor compatibility and computational efficiency. The main contributions of the paper are as follows:
  • We introduce novel 6G architectures and network simplification through user control plane separation and seamlessly integrating network domains through Agentic AI, enabling flexible, scalable service deployments while overcoming the limitations of current SBA.
@@
  • We discuss advances in constrained Agentic AI techniques applied to network operations, focusing on energy-efficient AI models, real-time learning optimization, and secure, adaptive AI-driven processes for improved network management.
@@
  • We explore innovative technologies such as real-time serverless computing for dynamic orchestration of functions, autonomous cognitive agents for decentralized network operations, goal-oriented communication protocols, cloudification of the RAN in space, and neural radio protocol stacks to support advanced, efficient and adaptive next-generation networks.
@@
The rest of the paper is organized as follows: Section II discusses new 6G architecture models that leverage Agentic AI for flexible service provisioning, efficient resource management and real-time learning under energy and security constraints. It highlights serverless computing frameworks, decentralized cognitive agents for adaptive network management, goal-oriented communication protocols for bandwidth and energy efficiency, cloudification of Radio Access Network (RAN) functions using space-based data centers, and customizable radio protocol stacks for real-time adaptation in 6G networks. Section III provides an experimental setup and an evaluation of agent-based AI within the Vehicle to Everything (V2X) scenario. Finally, Section IV summarizes the key innovations discussed and outlines the key steps to realizing the potential of 6G networks, focusing on scalability, intelligence and efficiency.
@@ II-A Novel Architectural Solutions for Simplified Networks with Agentic AI
Fig. 1 provides a layered diagram illustrating the evolution from 5G to 6G. A primary objective in the design of future 6G networks is the simplification of network architectures to facilitate more efficient service deployment and operation using AI agents. In this context, the separation of the User and Control planes allows for more flexible network management and resource allocation. This separation supports seamless integration of new network domains and ensures that devices can function as autonomous network nodes with the help of AI-support services. Such architectures allow for uninterrupted user roaming across various operators and network technologies in an adaptive manner. The limitations of the existing SBA in 5G networks are evident when considering the evolution toward 6G. Early models, such as the Hexa-X II framework [1], offer a foundation for addressing these shortcomings. However, further architectural refinements are needed to fully realize the potential of 6G, particularly with regard to supporting dynamic and adaptive services. Fig. 2 shows the network diagram that uses simplified user-control plane architecture with seamless domain integration using Agentic AI, where AI agents will be designed to optimize the services, communication, flow of data using context, semantics, and sustainability characteristics. Hence, realizing the true potential of 6G.
Figure 1: A layered diagram illustrating the evolution from 5G to 6G.
@@ II-C Real time serverless computing with Agentic AI for 6G Networks
Serverless computing, particularly in the form of FaaS [6], offers a promising solution for the dynamic orchestration of network services while employing Agentic AI as shown in Fig. 4. Achieving seamless scaling is difficult in existing network architectures due to the accessibility issues. By decoupling the execution of services from the underlying infrastructure, serverless computing enables the deployment of lightweight, highly-scalable functions across the network capable of automating and optimizing the task, including at the edge. In the context of 6G integrated with Agentic AI, real-time serverless computing can address several key challenges. These include mitigating the cold start problem, orchestrating high-dimensional tasks across multiple network domains, and ensuring security in distributed computing environments. The serverless computing with AI agents can also address the seamless scaling issues by employing proactive resource allocation and function pooling. The ability to dynamically allocate resources based on real-time network conditions is critical to achieving the low-latency, high-reliability performance required in next-generation networks. The latency of serverless computing (FaaS) with Agentic AI in real time can vary greatly depending on the platform, complexity of the function and network conditions. As detailed in [7], cold start latency can range from 100 milliseconds to several seconds. In a typical machine learning application for image classification, cold start latencies of about 2-3 seconds. Warm start latency, which occurs when an agent is called for a particular function shortly after a previous call, is typically in the range of 10-100 milliseconds and in a cloud scenario where a device sends image files to an AWS Lambda function deployed in a remote Amazon region, network latencies were between 70 and 120 milliseconds. In an edge scenario, where the device sends files from the same region as the Lambda function, network latencies were reduced to 0.5 to 2 milliseconds.
@@ II-D Autonomous Cognitive Agents in Network Architectures
The use of autonomous cognitive agents in network architectures represents a significant shift from the traditional, centralized network management models as shown in Fig. 5. Multi-Agent Systems (MAS), potentially using Large Language Models (LLMs), enable decentralized decision making where individual agents interact and cooperate to achieve network goals [16]. While it can be argued that the LLMs require computationally intensive operations, which could hinder the realization of the said network architecture, various techniques such as low-rank adapters (LoRA), quantization and others can be used to mitigate the computational complexity issues of the LLMs while reducing the computational complexity of the system. These agents can dynamically invoke each other, enabling a flexible and adaptable network infrastructure that operates in a more organic and decentralized manner. The design of architectures that support autonomous agents requires careful consideration of scalability, security and the seamless integration of service composition and knowledge handling. Decentralizing network operations not only improves efficiency, but also increases the network’s ability to adapt to changing conditions, providing robustness and resilience.
Figure 5: A decentralized agent-based network structure where autonomous agents communicate and coordinate without a central controller.
@@ II-E Neural Radio Protocols Stacks (NRPS) with Agentic AI
Finally, Table I provides a comparison of the 6G enablers using Agentic AI as discussed in this paper along with their features, advantages and disadvantages.
@@ III Experimental Setup and Evaluation
To validate Agentic AI for next-generation network architectures, we conduct a simple experiment to optimize packet delivery ratio and connection robustness in a V2X environment. For the development of AI agents, we use LLaMA2-13B and LLaMA2-70B as our foundational pre-trained models. We design three agents from the perspective of performance optimization, sustainability (reducing carbon footprint) and spectrum efficiency using crewAI [^3] https://github.com/crewAIInc/crewAI and develop a Retrieval Augmented Generation (RAG) using the Telecom Q&A [2], ArXiv papers on telecommunication/5G/6G and 3GPP documentation. Working together is at the heart of CrewAI. When multiple AI agents collaborate, they can tackle complex problems that would overwhelm a single agent working alone. By sharing what they know and dividing up tasks, the agents create a multiplier effect - their combined intelligence lets them solve challenges more effectively than any one agent could on its own. Think of it like a well-coordinated team where each member brings their unique strengths to the table. We used about 60 ArXiv papers for the development of RAG. To make it fair and simple, we did not apply any pre-processing or specialized curation of information to train AI agents, other than feeding the RAG. We modified the method proposed in [13] and performed an evaluation without the use of AI agents. We simulate the scenario described above for 100 - 500 devices. The mobility parameter was set to 10 - 30 m/s for vehicles in the V2X environment. Interference was set to ”Moderate to High” due to multiple active links, and the topology was set to random and clustered node placement. We evaluate performance based on the packet delivery ratio (PDR), i.e. the percentage of packets successfully delivered, and the robustness of the connection, i.e. the mean signal-to-noise ratio (SNR) and retransmission rate. Note that we repeat the simulation for a different number of parameters and give average results accordingly. Table II shows the comparative analysis for the parameter optimization of PDR and Link Robustness in V2X with and without Agentic AI. These results show that Agentic AI outperforms conventional methods when optimizing parameters, even with a simple RAG and few documents passed as training data. The Agentic AI approach achieves lower latency (20ms vs. 50ms), fewer retransmissions (17% vs. 5%), improved link robustness (20dB vs. 30dB), and higher packet delivery ratios (95% vs. 79%), respectively. We believe the improvement is due to the agent’s ability to dynamically adapt to changes in the network while optimizing resources and predicting failures in real time.
@@
[…]
IV Conclusion

### v3

@@ I Introduction
One of the key innovations of this paper lies in the design of simplified network architectures that separate the control and user planes, allowing operators to deploy and manage services more efficiently through AI Agents while seamlessly integrating new network domains. This approach also addresses the limitations of existing service-based architectures (SBA) and provides a clear path for the long-term evolution of 6G networks. The increased flexibility of such architectures ensures that devices can operate as dynamic network nodes, enabling uninterrupted roaming across multiple operators and network technologies. In addition to architectural simplification, the integration of AI Agents into network management [2] opens up new opportunities for optimization, especially in environments with energy constraints. Agentic AI-based networks can make intelligent decisions based on real-time data, but these models must work within strict energy and security constraints to ensure sustainable and secure operations. The concept of constrained AI operations, which is explored in this paper, is crucial for achieving these goals. It focuses on optimizing energy consumption and learning in real-time while maintaining security protocols. In addition, the emergence of serverless computing and function-as-a-service (FaaS) offers new solutions for the orchestration of services in distributed network environments. Serverless computing decouples application execution from the underlying infrastructure and enables real-time execution of lightweight functions at the network edge. This paradigm shift supports the flexible placement of AI agents and orchestration of tasks, a key requirement for the future 6G ecosystem where services must dynamically adapt to changing network conditions.
@@
Another fundamental component of future 6G architectures is the use of autonomous cognitive agents that introduce decentralized decision-making processes [3]. These agents work without a central controller and enable dynamic and spontaneous interactions that optimize network resources and operations. This is a significant departure from traditional static architectures and enables more adaptable and resilient networks that can meet future requirements. The concept of goal-oriented communication, a revolutionary approach that enables AI-equipped devices to transmit only the information necessary to achieve specific goals [4]. This reduces unnecessary communication, computing load and energy consumption, while allowing devices to communicate more efficiently. Such innovations are crucial for next-generation machine-to-machine (M2M) communication, where optimizing resource usage is key. In addition, the potential for cloudification and management of RAN functions in space underscores the increasing role of distributed data centers, including those in space, in the dynamic orchestration of RAN functions [5]. This approach contrasts with existing static methods of resource allocation and provides a more flexible and adaptable solution that can meet the diverse requirements of 6G communication scenarios. Cloudification of RAN functions is expected to play a critical role in achieving ultra- reliable, low-latency communications for a wide range of applications. The concept of Neural Radio Protocol Stacks (NRPS) can be used to automate the generation of radio protocol stacks tailored to specific network requirements using Agentic AI techniques. NRPSs can be a significant advance in the development of customizable and efficient communication protocols. By integrating Agentic AI into the development of radio stacks [6], these solutions can offer a high degree of flexibility and adaptability, addressing issues such as interpretability, cross-vendor compatibility and computational efficiency. The main contributions of the paper are as follows:
  • We propose a novel 6G architectural framework that decouples the user and control planes and leverages Agentic AI for seamless cross-domain integration, enabling scalable and adaptive service deployment beyond the limitations of current SBA.
@@
  • We introduce constrained Agentic AI techniques tailored for network operations, focusing on energy-efficient and secure real-time learning, with applications in dynamic resource allocation and threat-aware autonomous management.
@@
  • We explore the integration of real-time serverless computing with Agentic AI for agile function orchestration, autonomous cognitive agents for decentralized network control, space-cloudified RAN, and neural radio protocol stacks for adaptive and efficient next-generation systems.
  • We validate the effectiveness of Agentic AI through a V2X-based experimental study, demonstrating improvements in delivery packet ratio, latency, and link robustness compared to conventional methods.
@@
The rest of the paper is organized as follows: Section II discusses new 6G architecture models that leverage Agentic AI for flexible service provisioning, efficient resource management and real-time learning under energy and security constraints. It highlights serverless computing frameworks, decentralized cognitive agents for adaptive network management, goal-oriented communication protocols for bandwidth and energy efficiency, cloudification of radio access network (RAN) functions using space-based data centers, and customizable radio protocol stacks for real-time adaptation in 6G networks. Section III provides an experimental setup and an evaluation of agent-based AI within the vehicle-to-everything (V2X) scenario. Finally, Section V summarizes the key innovations discussed and outlines the key steps to realizing the potential of 6G networks, focusing on scalability, intelligence and efficiency.
@@ II-A Novel Architectural Solutions for Simplified Networks with Agentic AI
A primary objective in the design of future 6G networks is the simplification of network architectures to facilitate more efficient service deployment and operation using AI agents. The limitations of the existing SBA in 5G networks are evident when considering the evolution toward 6G. Early models, such as the Hexa-X II framework [7], offer a foundation for addressing these shortcomings. However, further architectural refinements are needed to fully realize the potential of 6G, particularly with regard to supporting dynamic and adaptive services. The separation of the user and control planes is a foundational step toward flexible 6G architectures. AI agents play a critical role in dynamically managing this separation by continuously monitoring traffic patterns, resource demands, and network states. For instance, control plane agents can autonomously adjust routing policies, session management rules, and QoS parameters based on real-time analytics, while user plane agents optimize packet forwarding paths and load balancing. This decoupling allows for independent scaling and updating of each plane. Moreover, AI agents enable intent-based orchestration across both planes, like translating high-level service requirements into low-level configuration actions, thus reducing manual intervention while enhancing operational agility. Through semantic understanding and context-aware decision making, these agents ensure that control and user plane functions remain synchronized while adapting to heterogeneous network conditions. Fig. 1 shows the network diagram that uses simplified user-control plane architecture with seamless domain integration using Agentic AI, where AI agents will be designed to optimize the services, communication, flow of data using context, semantics, and sustainability characteristics. Hence, realizing the true potential of 6G.
From a standardization perspective, the separation of control and user planes with Agentic AI introduces new requirements for interface definitions and protocol specifications. Current 3GPP standards for SBA would need extensions to support AI-agent interactions, particularly for intent-based networking interfaces and cross-domain coordination. Ongoing initiatives in ETSI ISG F5G [^1] https://www.etsi.org/technologies/fifth-generation-fixed-network-f5g and ITU-T FG-NET-2030222https://www.itu.int/en/ITU-T/focusgroups/net2030/Pages/default.aspx provide preliminary frameworks, but specific standards for AI-agent interoperability remain underdeveloped.
@@ II-C Real time serverless computing with Agentic AI for 6G Networks
Serverless computing, particularly in the form of FaaS [10], offers a promising solution for the dynamic orchestration of network services while employing Agentic AI as shown in Fig. 3. Achieving seamless scaling is difficult in existing network architectures due to the accessibility issues. By decoupling the execution of services from the underlying infrastructure, serverless computing enables the deployment of lightweight, highly-scalable functions across the network capable of automating and optimizing the task, including at the edge. In the context of 6G integrated with Agentic AI, real-time serverless computing can address several key challenges. These include mitigating the cold start problem, orchestrating high-dimensional tasks across multiple network domains, and ensuring security in distributed computing environments. The serverless computing with AI agents can also address the seamless scaling issues by employing proactive resource allocation and function pooling. The ability to dynamically allocate resources based on real-time network conditions is critical to achieving the low-latency, high-reliability performance required in next-generation networks. The latency of serverless computing FaaS with Agentic AI in real time can vary greatly depending on the platform, complexity of the function and network conditions. As detailed in [11], cold start latency can range from 100 milliseconds to several seconds. In a typical machine learning application for image classification, cold start latencies of about 2-3 seconds. Warm start latency, which occurs when an agent is called for a particular function shortly after a previous call, is typically in the range of 10-100 milliseconds and in a cloud scenario where a device sends image files to an AWS Lambda function deployed in a remote Amazon region, network latencies were between 70 and 120 milliseconds. In an edge scenario, where the device sends files from the same region as the Lambda function, network latencies were reduced to 0.5 to 2 milliseconds.
@@ II-D Autonomous Cognitive Agents in Network Architectures
The use of autonomous cognitive agents in network architectures represents a significant shift from the traditional, centralized network management models. Multi-Agent Systems (MAS), potentially using large language models (LLM), enable decentralized decision making where individual agents interact and cooperate to achieve network goals [12]. While it can be argued that the LLMs require computationally intensive operations, which could hinder the realization of the said network architecture, various techniques such as low-rank adapters (LoRA), quantization and others can be used to mitigate the computational complexity issues of the LLMs while reducing the computational complexity of the system. These agents can dynamically invoke each other, enabling a flexible and adaptable network infrastructure that operates in a more organic and decentralized manner. The design of architectures that support autonomous agents requires careful consideration of scalability, security and the seamless integration of service composition and knowledge handling. Decentralizing network operations not only improves efficiency, but also increases the network’s ability to adapt to changing conditions, providing robustness and resilience.
@@ II-E Neural Radio Protocols Stacks (NRPS) with Agentic AI
Lets consider the domain of channel state information (CSI) compression in order to discuss the practical applicability of NRPS. Traditional CSI feedback mechanisms are often inefficient under dynamic channel conditions. With NRPS, an AI agent can learn to compress CSI feedback adaptively by selecting only the most informative features based on real-time channel characteristics, thereby reducing overhead while maintaining link reliability. Similarly, in spectrum sharing scenario, NRPS-enabled agents can autonomously negotiate and allocate spectral resources among multiple operators by modeling spectrum usage patterns and interference constraints. For instance, in a shared license/unlicensed band scenario, agents can dynamically adjust transmission parameters such as power and modulation to maximize coexistence efficiency, leveraging real-time sensing and multi-agent collaboration. Finally, Table I provides a comparison of the 6G enablers using Agentic AI as discussed in this paper along with their features, advantages and disadvantages.
@@ III Experimental Setup and Evaluation
To validate Agentic AI for next-generation network architectures, we conduct a simple experiment to optimize packet delivery ratio and connection robustness in a V2X environment. The emulation of V2X environment in our experimental setup leveraged 100 to 500 devices moving at speeds between 10-30 m/s under moderate to high interference conditions. Node placement was randomized and clustered to simulate urban and highway scenarios. We implemented three specialized AI agents using LLaMA2-13B and LLaMA2-70B models, each optimized for performance, sustainability, and spectral efficiency. These agents were integrated via a crewaI framework [^5] https://github.com/crewAIInc/crewAI and supported by a Retrieval Augmented Generation (RAG) system built from 60 ArXiv papers, Telecom Q&A datasets [14], and 3GPP documentation. The V2X environment for vehicle mobility was simulated using simulation of urban mobility (SUMO) [^6] https://sumo.dlr.de/docs/Downloads.php and network emulation was simulated using NS-3. We used the scripts from the repository [^7] https://github.com/addola/NS3-HelperScripts/tree/master/examples/SUMOTraceExample for SUMOTrace, to gather the results. We used the socket connection to communicate between the Python agent code and the NS-3 simulation. The agents collaboratively adjusted transmission parameters such as power control, modulation schemes, and retransmission threshold in real time. When multiple AI agents collaborate, they can tackle complex problems that would overwhelm a single agent working alone. By sharing what they know and dividing up tasks, the agents create a multiplier effect, i.e. their combined intelligence lets them solve challenges more effectively than any one agent could on its own. Think of it like a well-coordinated team where each member brings their unique strengths to the table.
The experimental setup was designed to emulate 5G/6G usecase, i.e. URLLC for vehicular safety communications in V2X environment. The aforementioned setup assumes that the scenario models the transmission of cooperative awareness messages (CAMs) and decentralized environmental notification messages (DENMs) between vehicles and infrastructure, which are essential for collision avoidance and cooperative driving. The parameters are also configured according to 3GPP Release 16 and ETSI specifications for NR-V2X, including sub-6 GHz band operation, 10-100 MHz channel bandwidth, and slot-based transmission. The mobility patterns generated in SUMO reflect realistic urban scenarios with vehicle densities. The alignment of the experimental setup with standardized 5G/6G use case represents a real-world network deployment.
To make it fair and simple, we did not apply any pre-processing or specialized curation of information to train AI agents, other than feeding the RAG. We modified the method proposed in [15] and performed an evaluation without the use of AI agents. We evaluate performance based on the packet delivery ratio (PDR), i.e. the percentage of packets successfully delivered, and the robustness of the connection, i.e. the mean signal-to-noise ratio (SNR) and retransmission rate. Note that we repeat the simulation for a different number of parameters and give average results accordingly. Table II shows the comparative analysis for the parameter optimization of PDR and Link Robustness in V2X with and without Agentic AI. The results demonstrate that Agentic AI outperforms conventional methods across all metrics. The improvement in PDR (95% vs. 79%) and mean SNR (30dB vs. 20dB) highlights the agents’ ability to adapt to dynamic channel conditions. The reduction in retransmission rate (5% vs. 17%) and latency (20ms vs. 50ms) underscores the efficiency of AI-driven parameter tuning. These gains are attributed to the agents’ collaborative reasoning, which enables proactive resource allocation and failure prediction. The use of RAG further allowed the agents to leverage domain knowledge without extensive retraining, illustrating the practicality of lightweight AI integration into operational networks.
@@
Latency 50 ms ($\pm$ 10 ms) 20 ms ($\pm$ 5 ms)
IV Open Issues, Challenges, and Future Directions
Although the integration of Agentic AI into 6G networks presents several benefits and promises to overcome issues related to existing iterations of communication system, it has still unresolved challenges that require immediate research attention. We highlight some of open issues concerning Agentic AI based 6G networks and provide associated future directions to address those issues.
Real-time Orchestration of Serverless Functions: Current FaaS platforms struggle with the cold-start latencies exceeding 100ms, which is unacceptable for URLLC scenarios. This latency bottleneck is exacerbated by the stateless nature of serverless functions when maintaining AI agent context across multiple invocations. Future research should explore predictive function warming algorithms using temporal pattern analysis and stateful serverless containers that preserve agent memory across executions, potentially reducing cold-start times to under 10ms, especially for mission-critical applications.
Interpretability and Trustworthiness of NRPS: While AI-generated protocols can dynamically adapt to channel conditions, their black-box nature raises concerns about operational transparency and compliance with regulatory requirements. The inability to explain why specific protocol parameters were selected hinders adoption in safety-critical systems like autonomous vehicle communications. Future works should investigate hybrid symbolic-AI approaches that combine neural networks with human-curated RAG systems, creating auditable decision trails while maintaining adaptive capabilities. Developing standardized testing frameworks for NRPS verification across different vendor implementations will also be essential for commercial deployment.
Sustainability and Energy Usage for Distributed AI agents: The energy sustainability poses a fundamental constraint for widespread 6G deployment. As AI models grow in complexity, their computational demands conflict with the energy efficiency goals of green networking. Current LLMs require significant processing power, creating a carbon footprint that undermines network sustainability objectives. Research directions should focused on specialized TinyLLMs for edge devices, and renewable-energy-aware agent scheduling that aligns computational workloads with available green energy sources.
Multi-vendor Interoperability: Standardization gaps represent a critical barrier to multi-vendor interoperability in Agentic AI networks. The absence of unified interfaces for agent-to-agent communication and agent-to-network function interaction creates proprietary silos that limit scalability. Current 3GPP specifications lack definitive standards for AI agent authentication, service level agreements, and conflict resolution mechanisms. Future standardization efforts should prioritize open API specifications for agent communication, reference architectures for AI-native networks, and certification frameworks that ensure cross-vendor compatibility.
Security vulnerabilities in Multi-agent Systems: The security aspect require urgent attention to mitigate vulnerabilities in Agentic AI systems. The distributed nature of Agentic AI creates expanded attack surfaces, including model poisoning with malicious agents, adversarial attacks on decision processes, and privacy breaches through inference attacks. Traditional security mechanisms are insufficient for detecting sophisticated threats against collaborative AI systems. Future research should develop byzantine-resistant aggregation techniques for distributed learning, explainability-based anomaly detection that identifies aberrant agent behavior, and homomorphic encryption methods that allow secure computation on encrypted network data without compromising agent functionality.
Standardization of Agentic AI protocols: Standardization represents a critical pathway for the adoption of Agentic AI in 6G networks. Key standardization bodies including 3GPP, ITU-R, and IETF will need to address several dimensions, which include interface standardization between AI agents and network functions, certification frameworks for AI model safety and security in critical communications, performance benchmarking methodologies for AI-driven network operations, and ethical guidelines for autonomous network decision-making. The timeline for these standardization efforts aligns with 3GPP Release 19-20 timeframe, where initial studies on AI/ML for networks are already underway. Collaboration between academic researchers, industry stakeholders, and standard organization needs to be established in order to develop a coherent framework and comprehensive standards for Agentic AI network integration that enables innovation while ensuring interoperability and security.
V Conclusion